OpenStudy (anonymous):

solve the equation (x-4)(3x+2)=0

OpenStudy (austinl):

Ok, for this problem what would you do first?

OpenStudy (anonymous):

solve for x

OpenStudy (anonymous):

x=4,-2/3

OpenStudy (anonymous):

is it correct?

OpenStudy (austinl):

First I would FOIL out the left side of the problem.

OpenStudy (austinl):

What do you get?

OpenStudy (anonymous):

x-4=0 x=4 3x+2=0 3x=-2 x=-2/3

OpenStudy (austinl):

Seems good to me.